# Fait: Fault-Aware Fine-Tuning for Better Code Generation

### v1

@@ 3.3.2. Settings
All experiments are conducted on a machine with eight Tesla A800 GPUs, each with 80 GB of memory per GPU. is set to 2, which means the weight range of is . All models are trained for 1 epoch. Considering that Fait is designed to enable instruction-tuned LLMs to emphasize error-sensitive segments from their original instruction-tuning datasets, we apply relatively low learning rates during training. Specifically, we use a learning rate of 5e-6 with a linear scheduler and warm-up across all LLMs. The max sequence length is 1024. For inference evaluation, we use greedy decoding to ensure deterministic outputs, which also aligns with prior studies (Chen et al., 2021; Mu et al., 2024).
4. Results
@@
RQ2: How do different components of the Fait method contribute to LLMs’ performance?
RQ3: Does Fait demonstrate generalizability across different LLMs and their corresponding instruction-tuning datasets?
RQ4: Does Fait work for instruction-tuned LLMs whose instruction-tuning dataset is closed-source?
4.1. RQ1: Overall Effectiveness
In this RQ, we evaluate the effectiveness of our approach by applying it to several Instruction-tuned LLMs using their corresponding instruction-tuning datasets and assess their performance against three baselines:
Base Models: We use the original instruction-tuned LLMs without any additional Fait as our base models. This comparison demonstrates the absolute improvement achieved through Fait. Specifically, we select three representative instruction-tuned LLMs: MagiCoder-CL (Wei et al., 2024), MagiCoder-DS (Wei et al., 2024) and SemCoder-S (Ding et al., 2024) as our base models.
Closed-Source Models: We include GPT-3.5-Turbo (OpenAI, 2022) and GPT-4-Turbo (OpenAI and Josh Achiam, 2024) as the closed-source baseline. This comparison illustrates the performance gap between our fault-aware fine-tuned LLMs and advanced closed-source LLMs.
@@ 4.1. RQ1: Overall Effectiveness
We choose the instruction-tuning dataset evol-codealpaca-v1 (Luo et al., 2023) as our training dataset. This dataset is evolved from a seed dataset with GPT-4-Turbo, containing 110K high-quality data points. For MagiCoder-CL and MagiCoder-DS, this dataset is their original instruction-tuning dataset. For SemCoder-S, this dataset is a subset of its original instruction-tuned dataset, which is not fully open-sourced.
@@
Table 1 presents the performance of LLMs with Fait and the baselines across HumanEval(+), MBPP(+), and BigCodeBench. Overall, LLMs with Fait demonstrate substantial improvements in code generation. We observe that LLMs with Fait show average relative performance improvements of 4.8% over the base model and 4.9% over LLMs with Standard-SFT on HumanEval(+), MBPP(+), and BigCodeBench. Notably, with Fait, SemCoder-S with only 7B parameters outperforms the closed-source GPT-3.5-Turbo on HumanEval(+) and MBPP(+), and MagiCoder-DS outperforms GPT-3.5-Turbo on HumanEval(+). Both LLMs achieve comparable performance to GPT-3.5-Turbo on BigCodeBench, further validating the exceptional effectiveness of Fait in enhancing code generation capabilities. While the improvement on BigCodeBench-Full is modest, our approach shows more gains on BigCodeBench-Hard (e.g., 20.8% relative improvement for SemCoder-S). This is likely because more challenging problems contain more error-sensitive segments, and Fait is designed to guide LLMs to handle these error-sensitive segments, thus showing greater effectiveness on difficult programming tasks.
@@
When comparing the effects of Fait versus Standard-SFT on base models, we observe that Standard-SFT provides limited improvements and sometimes even weakens the base models. For example, SemCoder-S with Standard-SFT achieves only 0.8% relative performance improvement on HumanEval(+) and suffers 1.9% relative performance decline on MBPP+. This suggests that simply reinforcing the coarse-grained instruction-response mappings on their existing dataset provides minimal benefits, as these models have already captured these general mappings well during their initial instruction tuning. In contrast, Fait goes beyond simply continuing training on the original dataset, enabling LLMs to learn and memorize the mappings between problems and fine-grained error-sensitive segments in code. This targeted approach helps models distinguish between correct implementations and similar-looking but incorrect solutions, thereby generating accurate solutions when encountering target problems.
@@
To further figure out the reasons for Fait improving LLMs’ ability to generate functionally correct code, we manually inspect the test results. Based on our analysis, Fait demonstrates two main advantages over both the original model and Standard-SFT:
@@
First, Fait can learn diverse error-sensitive segments to better recognize and focus on implementation details that are prone to errors, while the original model and Standard-SFT only learn the overall mapping from problem to solution. This method improves the LLM’s attention to key implementation choices. For example, Figure 4 presents a comparison of the results of three versions of SemCoder-S on the HumanEval/135 task. In this example, an error-sensitive segment involves deciding whether to return the index of the target element itself i or the index of its previous element i-1. This distinction directly impacts the functional correctness of the implementation. Both the original model and the model with Standard-SFT incorrectly return the index of the previous element, while the model with Fait correctly returns the index of the target element. We manually examine the tasks from HumanEval(+) that are correctly solved after applying Fait but initially incorrect with the base models. We find that among these tasks, 63.6% of SemCoder-S’s improvements result from properly handling error-sensitive segments, with similar rates observed in MagiCoder-DS (71.4%) and MagiCoder-CL (72.7%). These results further confirm our method’s effectiveness in guiding LLMs to recognize error-sensitive segments.
@@
Second, by developing a deeper understanding of critical code segments, Fait also enhances overall code generation capabilities. By strategically emphasizing error-sensitive segments while maintaining appropriate weight for contextual elements, the LLM learns to identify and handle the crucial parts of implementations that determine correctness. Figure 5 demonstrates this using an example from SemCoder-S on HumanEval/110, which requires determining whether swapping elements between two lists can make all elements in lst1 even. This case highlights improvements that go beyond addressing specific error-sensitive segments. Both the original model and Standard-SFT fail to implement the correct verification logic to determine whether lst2 contains enough even numbers to replace odd numbers in lst1. In contrast, the Fault-Aware Fine-Tuned model correctly implements this logic, demonstrating enhanced general coding abilities rather than just handling error-sensitive parts.
4.2. RQ2: Component Analysis
@@
+Fait (Line Level) | 75.6 (72.0) | 78.8 (68.3) | 47.1 | 14.2 | ||||
+Fait (Token Level) | 76.2 (72.6) | 79.1 (68.5) | 47.3 | 14.2 | ||||
+Fait | 77.4 (74.3) | 79.6 (69.0) | 48.2 | 15.5 | ||||
@@
+Fait (Line Level) | 80.5 (76.8) | 83.1 (70.1) | 49.1 | 18.2 | ||||
+Fait (Token Level) | 81.1 (76.8) | 82.8 (70.1) | 48.3 | 16.9 | ||||
+Fait | 83.5 (78.7) | 83.1 (70.6) | 48.9 | 20.3 |
To understand how different components contribute to the effectiveness of Fait, we conduct ablation studies focusing on the impacts of multi-granularity and the loss function in this RQ.
Impact of Difference Granularity. We explore the impact of code difference granularity, which involves synthesizing line-level and token-level code differences to identify error-sensitive segments. Specifically, we conduct ablation experiments using MagiCoder-DS and SemCoder-S as base models and perform evaluation on three selected benchmarks. Table 2 shows the impact of different granularities of differences on Fait. We can observe that the combination of line-level granularity and token-level granularity yields maximum performance gains. For example, when applied to SemCoder-S, this approach achieves a relative average improvement of 4.9% across all benchmarks compared to the base model, compared to just 2.9% for line-level only and 2.4% for token-level only. This demonstrates that multi-granularity differences enable better error-sensitive segments, thereby enhancing LLMs’ code generation capabilities.
Impact of the Loss Function. To validate the effectiveness of our dynamic loss weighting design, we compare our dynamic weighting approach against a fixed weighting strategy where all error-sensitive tokens receive the same constant weight during training. We experiment with fixed weights in the range of on HumanEval(+). Due to the evaluation time constraints, we select SemCoder-S as the representative LLM for this ablation study, as it demonstrates the best overall performance with Fait. Notably, when the fixed weight equals 1, this configuration is equivalent to standard SFT. Figure 6 shows the performance of SemCoder-S with different fixed weights compared to our dynamic weighting approach. Our experimental results reveal two important findings: (1) A fixed weight of 2 yields better performance than other fixed weight values, suggesting that an appropriate constant weight can help the LLM recognize error-sensitive segments and enhance code generation capabilities. (2) Our dynamic weighting approach still outperforms the best fixed weighting configuration. This confirms that dynamically adjusting weights based on the LLM’s current discrimination ability provides better guidance for the LLM to focus on critical implementation details that differentiate correct solutions from their erroneous variants.
4.3. RQ3: The Generalization Capabilities
@@
+Fait | 73.7 (69.5) | 81.0 (67.2) | 47.9 | 21.6 |
In this RQ, we aim to explore the generalizability of Fait across different instruction-tuned LLMs when using their own instruction-tuning datasets. Specifically, we select two representative LLMs and their corresponding instruction datasets. 1) We select MagiCoder-DS and its corresponding dataset OSS-Instruct. This dataset is generated from open-sourced code by GPT-3.5-Turbo, and contains 75K samples. 2) SemCoder and its corresponding dataset PYX. This dataset consists of 95K samples, including comprehensive reasoning texts with executable code samples. The dataset is constructed with problem descriptions generated by GPT-3.5-Turbo and corresponding responses generated by GPT-4o-mini (OpenAI, 2024), creating high-quality instruction-response pairs with detailed reasoning. For each LLM, we process its corresponding dataset through our pipeline and evaluate performance on the same benchmarks used in RQ1. We select Base Models and Standard-SFT Models as our baselines. By using LLMs with different training paradigms and datasets with varying characteristics (open-sourced code versus detailed reasoning with executable samples), we can verify that our approach is not tied to specific LLM series or dataset properties, but rather provides universal benefits.
Table 3 shows the performance of LLMs on Humaneval(+), MBPP(+), and BigCodeBench after Fait and Standard-SFT. We can observe that Fait demonstrates robust generalization capabilities to different instruction-tuning LLMs and their corresponding datasets. For MagiCoder-DS and SemCoder, after Fait, the average performances across all benchmarks show relative improvements of 5.3% and 3.8% compared to the base models. In contrast, standard SFT yielded modest relative improvements of 1.4% for MagiCoder-DS and decreased performance by 2.1% for SemCoder. These results show that Fait’s benefits are not tied to specific dataset characteristics or model series. Instead, the approach effectively enhances diverse instruction-tuned LLMs by teaching them to focus on error-sensitive segments in correct solutions.
4.4. RQ4: Effectiveness on LLMs with Closed-Source Instruction Data
@@
+Fait | 81.7 (76.2) | 78.6 (66.9) | 44.3 | 16.9 |
A key question for the broader adoption is whether Fait can enhance LLMs whose original instruction-tuning datasets are not publicly available. To investigate this, we applied our method to two widely-used LLMs with closed-source training data in this RQ. Specifically, we choose CodeLlama-7B-Instruct (Roziere et al., 2023) and DeepSeek-
Coder-6.7B-Instruct (Guo et al., 2024) as base models. These LLMs are instruction-tuned on substantial but proprietary datasets - CodeLlama-7B-Instruct underwent instruction tuning on approximately 5B tokens of instruction data, while DeepSeekCoder-6.7B-Instruct is tuned on around 2B tokens. To test our approach without access to these original datasets, we select evol-codealpaca-v1 used in the main experiment as the training dataset. We select Base Models and Base Models with Standard SFT as our baselines and evaluate on Humaneval(+), MBPP(+) and BigCodeBench(+).
Table 4 shows the performance of these two LLMs using Standard-SFT and fault-fine-tuning on Humaneval(+), MBPP(+), and BigCodeBench. We can find that Fait is also applicable to LLMs with closed-source datasets. For CodeLlama-Instruct and DeepseekCoder-Instruct, after Fait, the average relative improvements across all benchmarks are 19.1% and 5.9%. By comparison, the standard SFT yield gains of 7.5% and 0.5%, respectively. This further demonstrates that fault-fine-tuning is also applicable to LLMs with closed-source datasets, showing strong applicability.
5. Discussion
Impact of in Loss. The parameter in our loss function controls the emphasis placed on error-sensitive tokens. We further experimentally investigate the performance impact of changing . Specifically, we conduct experiments with on Humaneval(+) and observe the performance changes of LLMs. Due to the evaluation time constraints, we select SemCoder-S as the representative LLM for this ablation study, as it demonstrates the best overall performance with Fait. Figure 7 illustrates the performance trends as varies. We can observe that provides an optimal balance between emphasizing error-sensitive tokens and maintaining attention to shared tokens. Additionally, it can be observed that in all cases, after Fait, the LLM’s performance matches or exceeds that of Standard-SFT, demonstrating the robustness to hyperparameter choices.

### v2

@@ III-B2 Settings
All experiments are conducted on a machine with eight Tesla A800 GPUs, each with 80 GB of memory. is set to 2, which means the weight range of is (1,2). Given that FGit is designed for already instruction-tuned LLMs, where the objective is not to train them extensively from a base state but rather to refine their existing capabilities to better handle error-sensitive segments present, we train all models for one epoch with a relatively low learning rate of 5e-6, aiming to gently adapt the model and prevent catastrophic forgetting of previously learned knowledge, while efficiently instilling the new focus on these critical segments [33, 34]. The max sequence length is 1024. For inference evaluation, we use greedy decoding to ensure deterministic outputs, which also aligns with prior studies [6, 26].
IV Results
@@
RQ2: How do different components of the FGit method contribute to LLMs’ performance?
RQ3: Does FGit demonstrate generalizability across different LLMs and their corresponding instruction-tuning datasets?
RQ4: Does FGit work for instruction-tuned LLMs whose instruction-tuning dataset is closed-source?
IV-A RQ1: Overall Effectiveness
In this RQ, we evaluate the effectiveness of our approach by applying it to several Instruction-tuned LLMs using their corresponding instruction-tuning datasets and assess their performance against four baselines:
Base Models: We use the original instruction-tuned LLMs without any additional FGit as our base models. This comparison demonstrates the absolute improvement achieved through FGit. Specifically, we select three representative instruction-tuned LLMs: MagiCoder-CL [2], MagiCoder-DS [2] and SemCoder-S [18] as our base models.
Closed-Source Model: We include GPT-3.5-Turbo [19] as the closed-source baseline to illustrate the performance gap between our fault-guided fine-tuned LLMs and the advanced closed-source LLM.
@@ IV-A RQ1: Overall Effectiveness
Other Open Source Instruction-Tuned Code LLMs: We include other instruction-tuned versions of the same base models, which are trained on additional data generated by stronger LLMs (e.g., GPT-4) in addition to Evol-Instruct. Specifically, we include WaveCoder-Ultra [5], OpenCodeInterpreter[3],
and AlchemistCoder [35].
This comparison aims to illustrate how the performance of the selected LLMs with FGit compares to that of current popular instruction-tuned models based on the same foundational model.
We choose Evol-instruct [2], a decontaminated version of evol-codealpaca-v1 [36], which contains numerous high-quality instruction-following data, as our training dataset. This dataset is decontaminated by removing data that contain docstrings or solutions from multiple benchmarks [6, 15, 37, 38]. We conduct an additional decontamination of this dataset for docstrings or solutions from BigCodeBench following [2], and find no overlap. For MagiCoder-CL and MagiCoder-DS, this dataset is their original instruction-tuning dataset. For SemCoder-S, this dataset is a subset of its original instruction-tuned dataset, which is not fully open-sourced.
@@
We report results consistently from the EvalPlus leaderboard111https://evalplus.github.io/leaderboard.html and BigCodeBench leaderboard222https://bigcode-bench.github.io. Table I presents the performance of LLMs with FGit and the baselines across HumanEval(+), MBPP(+), and BigCodeBench. Overall, LLMs with FGit demonstrate substantial improvements in code generation. We observe that LLMs with FGit show average relative performance improvements of 4.8% over the base model and 4.9% over LLMs with Standard-SFT on HumanEval(+), MBPP(+), and BigCodeBench. Notably, with FGit, SemCoder-S with only 7B parameters outperforms GPT-3.5-Turbo on HumanEval(+) and MBPP(+), and MagiCoder-DS outperforms GPT-3.5-Turbo on HumanEval(+). Both LLMs achieve comparable performance to GPT-3.5-Turbo on BigCodeBench, further validating the exceptional effectiveness of FGit in enhancing code generation capabilities. While the improvement on BigCodeBench-Full is modest, our approach shows more gains on BigCodeBench-Hard (e.g., 20.8% relative improvement for SemCoder-S). This is likely because more challenging problems contain more error-sensitive segments, and FGit is designed to guide LLMs to handle these error-sensitive segments, thus showing greater effectiveness on difficult programming tasks.
@@
When comparing the effects of FGit versus Standard-SFT on base models, we observe that Standard-SFT provides limited improvements and sometimes even weakens the base models. For example, SemCoder-S with Standard-SFT achieves only 0.8% relative performance improvement on HumanEval(+) and suffers 1.9% relative performance decline on MBPP+. This suggests that simply reinforcing the coarse-grained instruction-response mappings on their existing dataset provides minimal benefits, as these models have already captured these general mappings well during their initial instruction tuning. When compared with other instruction-tuned models, we observe that by enabling models to further learn the error-sensitive segments within the original dataset through FGit, their performance can achieve comparable or even superior results.
@@
To further figure out the reasons for FGit improving LLMs’ ability to generate functionally correct code, we manually inspect the results. Based on our analysis, FGit demonstrates two main advantages over both the original model and Standard-SFT:
@@
First, FGit can learn diverse error-sensitive segments to better recognize and focus on implementation details that are prone to errors, while the original model and Standard-SFT only learn the overall mapping from problem to solution. This method improves the LLM’s attention to key implementation choices. We manually examine the tasks from HumanEval(+) that are correctly solved by applying FGit but are incorrect when applying Standard-SFT. We find that for SemCoder-S, FGit correct 11 tasks where Standard-SFT failed. In 81.8% of these instances, Standard-SFT’s failure stemmed from incorrectly handling error-sensitive segments; these errors typically required modifications of three lines of code or fewer to be rectified. FGit, in contrast, provides a correct implementation, with similar rates observed in MagiCoder-DS (71.4%) and MagiCoder-CL (66.7%). For example, Figure 4 presents a comparison of the results of three versions of SemCoder-S on the HumanEval/135 task. In this example, an error-sensitive segment involves deciding whether to return the index of the target element itself i or the index of its previous element i-1. This distinction directly impacts the functional correctness of the implementation. Both the original model and the model with Standard-SFT incorrectly return the index of the previous element, while the model with FGit correctly returns the index of the target element. These results further confirm our method’s effectiveness in guiding LLMs to recognize error-sensitive segments.
@@
Second, by developing a deeper understanding of critical code segments, FGit also enhances overall code generation capabilities. By strategically emphasizing error-sensitive segments while maintaining appropriate weight for contextual elements, the LLM learns to identify and handle the crucial parts of implementations that determine correctness. Figure 5 demonstrates this using an example from SemCoder-S on HumanEval/110, which requires determining whether swapping elements between two lists can make all elements in lst1 even. This case highlights improvements that go beyond addressing specific error-sensitive segments. Both the original model and Standard-SFT fail to implement the correct verification logic to determine whether lst2 contains enough even numbers to replace odd numbers in lst1. In contrast, the Fault-Guided Fine-Tuned model correctly implements this logic, demonstrating enhanced general coding abilities rather than just handling error-sensitive parts.
IV-B RQ2: Component Analysis
@@
+FGit (Line Level) | 75.6 (72.0) | 78.8 (68.3) | 47.1 | 14.2 | ||||
+FGit (Token Level) | 76.2 (72.6) | 79.1 (68.5) | 47.3 | 14.2 | ||||
+FGit | 77.4 (74.3) | 79.6 (69.0) | 48.2 | 15.5 | ||||
@@
+FGit (Line Level) | 80.5 (76.8) | 83.1 (70.1) | 49.1 | 18.2 | ||||
+FGit (Token Level) | 81.1 (76.8) | 82.8 (70.1) | 48.3 | 16.9 | ||||
+FGit | 83.5 (78.7) | 83.1 (70.6) | 48.9 | 20.3 |
To understand how different components contribute to the effectiveness of FGit, we conduct ablation studies focusing on the impacts of multi-granularity and the loss function.
Impact of Difference Granularity. We explore the impact of code difference granularity, which involves synthesizing line-level and token-level code differences to identify error-sensitive segments. Specifically, we conduct ablation experiments using MagiCoder-DS and SemCoder-S as base models and perform evaluation on three selected benchmarks. Table II shows the impact of different granularities of differences on FGit. We can observe that the combination of line-level granularity and token-level granularity yields maximum performance gains. For example, when applied to SemCoder-S, this approach achieves a relative average improvement of 4.9% across all benchmarks compared to the base model, compared to just 2.9% for line-level only and 2.4% for token-level only.
We further investigate models’ ability to handle error-sensitive segments after combining both granularities, compared to the base models. We select the tasks from HumanEval(+) that are correctly solved after applying FGit but initially incorrect with the base models. We find that among these tasks, 63.6% of SemCoder-S’s improvements result from properly handling error-sensitive segments, with these specific error-sensitive segments-related corrections involving modifications of three lines or fewer. Similar rates are observed in MagiCoder-DS (71.4%). This demonstrates that multi-granularity differences enable better error-sensitive segments, thereby enhancing LLMs’ code generation capabilities.
Impact of the Loss Function. To validate the effectiveness of our dynamic loss weighting design, we compare our dynamic weighting approach against a fixed weighting strategy where all error-sensitive tokens receive the same constant weight during training. We experiment with fixed weights in the range of on HumanEval(+). Due to the evaluation time constraints, we select SemCoder-S as the representative LLM for this ablation study, as it demonstrates the best overall performance with FGit. Notably, when the fixed weight equals 1, this configuration is equivalent to standard SFT. Figure 6 shows the performance of SemCoder-S with different fixed weights compared to our dynamic weighting approach. Our experimental results reveal two important findings: (1) A fixed weight of 2 yields better performance than other fixed weight values, suggesting that an appropriate constant weight can help the LLM recognize error-sensitive segments and enhance code generation capabilities. (2) Our dynamic weighting approach still outperforms the best fixed weighting configuration. This confirms that dynamically adjusting weights based on the LLM’s current discrimination ability provides better guidance for the LLM to focus on critical implementation details that differentiate correct solutions from their erroneous variants.
IV-C RQ3: The Generalization Capabilities
@@
+FGit | 73.7 (69.5) | 81.0 (67.2) | 47.9 | 21.6 |
In this RQ, we aim to explore the generalizability of FGit across different instruction-tuned LLMs when using their own instruction-tuning datasets. Specifically, we select two representative LLMs and their corresponding instruction datasets. 1) We select MagiCoder-DS and its corresponding dataset OSS-Instruct. This dataset is generated from open-sourced code by GPT-3.5-Turbo, and contains 75K samples. 2) SemCoder and its corresponding dataset PYX. This dataset consists of 95K samples, including comprehensive reasoning texts with executable code samples. The dataset is constructed with problem descriptions generated by GPT-3.5-Turbo and corresponding responses generated by GPT-4o-mini [39], creating high-quality instruction-response pairs with detailed reasoning. For each LLM, we process its corresponding dataset through our pipeline and evaluate performance on the same benchmarks used in RQ1. To be noted, we also decontaminate the datasets, adhering to [2], and find no data overlap with our selected evaluation benchmarks. We select Base Models and Standard-SFT Models as our baselines. By using LLMs with different training paradigms and datasets with varying characteristics (open-sourced code versus detailed reasoning with executable samples), we can verify that our approach is not tied to specific LLM series or dataset properties, but rather provides universal benefits.
Table III shows the performance of LLMs on HumanEval(+), MBPP(+), and BigCodeBench after FGit and Standard-SFT. We can observe that FGit demonstrates robust generalization capabilities to different instruction-tuning LLMs and their corresponding datasets. For MagiCoder-DS and SemCoder, after FGit, the average performances across all benchmarks show relative improvements of 5.3% and 3.8% compared to the base models. In contrast, standard SFT yielded modest relative improvements of 1.4% for MagiCoder-DS and decreased performance by 2.1% for SemCoder. These results show that FGit’s benefits are not tied to specific dataset characteristics or model series. Instead, the approach effectively enhances diverse instruction-tuned LLMs by teaching them to focus on error-sensitive segments in correct solutions.
IV-D RQ4: Effectiveness on LLMs with Closed-Source Instruction Data
@@
+FGit | 81.7 (76.2) | 78.6 (66.9) | 44.3 | 16.9 |
A key question for the broader adoption is whether FGit can enhance LLMs whose original instruction-tuning datasets are not publicly available. To investigate this, we applied our method to two widely-used LLMs with closed-source training data in this RQ. Specifically, we choose CodeLlama-7B-Instruct [40] and DeepSeek-
Coder-6.7B-Instruct [41] as base models. These LLMs are instruction-tuned on substantial but proprietary datasets - CodeLlama-7B-Instruct underwent instruction tuning on approximately 5B tokens of instruction data, while DeepSeekCoder-6.7B-Instruct is tuned on around 2B tokens. To test our approach without access to these original datasets, we select Evol-Instruct, used in the main experiment, as the training dataset. We select Base Models, and Base Models with Standard SFT as our baselines and evaluate on HumanEval(+), MBPP(+) and BigCodeBench(+).
Table IV shows the performance of these two LLMs using Standard-SFT and fault-fine-tuning on HumanEval(+), MBPP(+), and BigCodeBench. We can find that FGit is also applicable to LLMs with closed-source datasets. For CodeLlama-Instruct and DeepseekCoder-Instruct, after FGit, the average relative improvements across all benchmarks are 19.1% and 5.9%. By comparison, the standard SFT yield gains of 7.5% and 0.5%, respectively. This further demonstrates that fault-fine-tuning is also applicable to LLMs with closed-source datasets, showing strong applicability.
V Discussion
Impact of hyperparameters. We first analyse the impact of the parameter in our loss function, which controls the emphasis placed on error-sensitive tokens, by changing . Specifically, we conduct experiments with on HumanEval(+) and observe the performance changes of LLMs. Due to the evaluation time constraints, we select SemCoder-S as the representative LLM for this ablation study, as it demonstrates the best overall performance with FGit. Figure 7(a) illustrates the performance trends as varies. We can observe that provides an optimal balance between emphasizing error-sensitive tokens and maintaining attention to shared tokens. Additionally, it can be observed that in all cases, after FGit, the LLM’s performance matches or exceeds that of Standard-SFT, demonstrating the robustness to hyperparameter choices.
Then we investigate the impact of the number of training epochs, which also influences the learning emphasis on error-sensitive segments. Specifically, we conduct experiments with epochs in using SemCoder-S and observe the performance changes on HumanEval(+). Figure 7(b) illustrates the performance trends as the number of training epochs varies. We find that an excessive number of training epochs can lead to a decline in performance. This might be because the model over-focuses on error-sensitive segments, neglecting the importance of surrounding tokens.